If the first and third of three consecutive odd integers are added, the result is 69 less than five times the second integer. Find all three integers.
Answer by ewatrrr(24785)


Hi
Let x, (x+2) and (x+4) represent the consecutive odd Integers respectively
Question states***
x + (x+4) = 5(x+2) - 69
Solving for x
2x + 4 = 5x - 59
63 = 3x
21 = x The consecutive odd Integers are 21,23,25
CHECKING our Answer***
46 = 5*23 - 69 = 115-69 = 46